# Cosmological constraints from $N$-body simulations for future spectroscopic galaxy surveys at $2\leq z\leq 3$

### v1

@@ Abstract
I Introduction
Cosmological models have been tested via various observations including cosmic microwave background (CMB) Hinshaw et al. (2013); Planck Collaboration et al. (2016), gravitational lensing Abbott et al. (2018); Hikage et al. (2019), galaxy redshift surveys Ivanov et al. (2020); d’Amico et al. (2020); Kobayashi et al. (2022), Type-Ia supernovae Suzuki et al. (2012); DES Collaboration et al. (2024), etc. Cold dark matter with a cosmological constant, the $\Lambda$CDM model, has successfully explained all of these observations so far. However, there exist discrepancies between some parameters determined from the early universe and those from the local universe, known as the Hubble tension (Verde et al., 2019) and the $S_{8}$ tension Abdalla et al. (2022). The Dark Energy Spectroscopic Instrument (DESI) DESI Collaboration et al. (2024), the latest galaxy survey, has reported cosmological constraints that support dynamical dark energy with $w>-1$. However, combining them with other datasets still includes the cosmological constant ($w=-1$) within $2\sigma$ in a large cosmological parameter space (e.g., Choudhury and Okumura (2024)). Although the geometry of the universe has been found flat by the CMB, the flatness of the universe has not been constrained in the late universe observations (Takada and Doré, 2015; Terasawa et al., 2022, 2024). One way to tackle these issues is to extend the observations of cosmological objects to the more distant universe Okumura et al. (2016). In addition, cosmological spectroscopic experiments at $z\gtrsim 2$ have the potential to put meaningful constraints on curvature (Takada and Doré, 2015), neutrino mass (Takada et al., 2006; Dvorkin et al., 2019), inflation (Takada et al., 2006), and dark energy (Ferraro and Wilson, 2019).
@@ I Introduction
II $N$-body Simulations and halo catalogs
We use $N$-body simulations run as part of the DARK QUEST project (Nishimichi et al., 2019). We employ $2048^{3}$ dark matter particles of mass $m_{p}=8.15875\times 10^{10}h^{-1}M_{\odot}$ in a cubic box on the side $2h^{-1}{\rm Gpc}$. We have the data set from three independent realizations and we specifically analyze clustering of dark matter halos of the snapshots at $z=2$ and $z=3$. Halos are identified using the ROCKSTAR algorithm (Behroozi et al., 2013). Their velocities and positions are determined by the average of the member particles within the innermost 10% of the subhalo radius. The halo mass, $M_{\rm h}$, is defined by a sphere with radius within which the enclosed average density is 200 times the mean matter density.
@@ II $N$-body Simulations and halo catalogs
III Redshift-space correlation functions
III.1 Measurements from $N$-body simulations
@@
×$\qquad\qquad\qquad\quad\times\left[\xi_{\ell,k}(r_{j})-\bar{\xi}_{\ell}(r_{j})\right],$ (2)
where $\bar{\xi}_{\ell}(r_{i})=N_{\rm mock}^{-1}\sum_{k=1}^{N_{\rm mock}}\xi_{\ell,k}(r_{i})$. We create fifty realizations from each simulation. Furthermore, in measuring the redshift-space density field, we regard each direction along the three axes of simulation boxes as the line of sight. Thus, we have $N_{\rm mock}=50\times 3\times 3=450$ realizations to construct the covariance matrix.
Figure 1: Multipoles of redshift-space correlation functions of halos at $z=2$ (left) and $z=3$ (right). The upper and lower rows show the results for the low- and high-mass halos, respectively. The red and blue points are the measured monopole and quadrupole moments from $N$-body simulations. The dashed curves are the model with the input cosmological parameters. The dotted curves are the best-fitting model at $60\leq r\leq 140{\hbox{$~{}h^{-1}$}{\rm~{}Mpc}}$ (see Sec. IV.2). Errors are the squared diagonal components of the covariance matrices.
@@
III.2 Linear theory
Here we present a theoretical model of the measured correlation functions to place cosmological constraints. Since theoretical models are derived in Fourier space, we first consider a model for the redshift-space power spectrum, $P(\mathbf{k})=P(k,\mu_{\bm{k}})$, with $k=|\bm{k}|$ and $\mu_{\bm{k}}$ the directional cosine. Given the underlying matter power spectrum, cosmological information is extracted from the dynamical and geometric distortions, which are probed, respectively, via redshift-space distortion (RSD) and Alcock-Paczynski (AP) effects (Alcock and Paczynski, 1979; Kaiser, 1987; Seo and Eisenstein, 2003; Matsubara, 2004; Okumura et al., 2008).
In this paper, we use the simplest model for RSD, the linear RSD model with a constant bias $b$, known as the Kaiser model (Kaiser, 1987). It is given by
@@ III.2 Linear theory
These multipole power spectrum can be converted into multipole correlation functions by
@@
To measure the correlation function, the halo position sampled with redshift and angular position need to be converted into the comoving positions by introducing a reference cosmology. The apparent mismatch between the reference and true cosmology causes a geometric distortion in the measured correlation function, known as the AP effect Alcock and Paczynski (1979). Taking this effect into account, the correlation functions presented above are modulated as (Xu et al., 2013)
@@
where $H(z)$ and $D_{\rm A}(z)$ are the expansion rate and angular diameter distance, respectively, and $r_{\rm s}$ the sound horizon scale at recombination time. Quantities with the subscript $\rm{fid}$ are computed using fiducial cosmological parameters (see Sec. I). Detecting anisotropies of BAO enables us to separate the AP effect from the dynamical distortion effect.
@@
IV Cosmological analysis and results
In this section, we investigate how well one can constrain cosmological models at redshifts $z>2$ using our high-redshift simulations. In Sec. IV.2, we present constraints on the growth rate and expansion rate using dynamical and geometric distortions. In Sec. IV.3, we constrain cosmological parameters by taking into account the cosmological dependence of the underlying power spectrum shape.
IV.1 Setup for likelihood analysis
We perform the likelihood analysis using the measured correlation functions of halos with the theoretical models described in Sec. III.2. The $\chi^{2}$ statistic is given by
@@ IV.1 Setup for likelihood analysis
IV.2 Dynamical and geometric constraints
@@
IV.3 Full-shape constraints
Here we utilize the cosmological information encoded in the underlying matter power spectrum and directly constrain cosmological parameters. We consider a nonflat $\Lambda$CDM model with six parameters in total, $\theta=(b,~{}\Omega_{\rm m},~{}\Omega_{\rm k},~{}H_{0},~{}A_{\rm s},~{}k_{*})$, where $A_{\rm s}$ is the amplitude of the primordial curvature perturbations. We fix $\Omega_{\rm b}$ to the fiducial value because the baryon density is tightly constrained by CMB and big-bang nucleosynthesis Aver et al. (2013); Cooke et al. (2014); Planck Collaboration et al. (2016). We further fix the dark energy equation-of-state to a cosmological constant, $w_{0}=-1$, as the high-$z$ clustering is not sensitive to $w$. On the other hand, the high-redshift universe is a useful laboratory to test the flatness of the universe, $\Omega_{\rm k}$ (Takada and Doré, 2015).
@@ IV.3 Full-shape constraints
Fig. 4 shows the two-dimensional constraints on pairs of cosmological parameters. Table 3 summarizes the one-dimensional marginalized constraints. The best-fitting value of $b$ obtained here is fully consistent with that from the dynamical and geometric constraints. We obtain $1.3\%$ and $2.0\%$ constraints on $\Omega_{\rm m}$ from the clustering of halos at $z=2$ and $z=3$, respectively. Similarly, we achieve constraints of $1.7\%$ and $2.4\%$ in $H_{0}$. The reported discrepancy between the values of $H_{0}$ measured using early and late universe probes is about $5-10\%$. Our result thus demonstrates that achieving the number density of $\bar{n}_{\rm g}\simeq 10^{-4}({\hbox{$~{}h^{3}{\rm~{}Mpc}^{-3}$}})$ over the large volume at $z\geq 2$ provides a powerful test of the Hubble tension, independent of the local universe ($z\simeq 0$) and early universe ($z\simeq 1100)$ measurements.
Note that here our analyses show certain offsets for $\Omega_{\rm m}$ and $H_{0}$, while the results shown in the previous subsection are consistent with the input within $1\sigma$ level for these samples. These shifts are driven by the inaccuracy of our linear theory model with the simple BAO damping on small scales, where cosmological information can still be extracted. In fact, Fig. 1 shows that the best-fit model fails to capture behavior on smaller scales than the BAO scale, especially at $z=2$, while it agrees with the measurements up to the BAO scale. This issue can be easily restored once one employs a more sophisticated model and our main conclusion of constraining cosmology from high redshifts is unchanged.
@@
V Summary and discussion
In this paper, we have studied the cosmological constraints expected from upcoming high-redshift galaxy surveys utilizing high-resolution $N$-body simulations at $z=2$ and 3. To obtain general conclusions, we did not consider specific galaxy populations or survey geometries but rather split our halo sample into low-mass and high-mass subsamples in a cubic box of side $2{\hbox{$~{}h^{-1}$}{\rm~{}Gpc}}$. We chose a conservative fitting range where a simple linear theory prediction recovered the input cosmological models. We considered two analysis methods, a template fitting for dynamical and geometric constraints and a full-shape analysis. With the former, we obtained $\sim 10\%$ constraint on $f(z)$ with the massive halo sample with $\bar{n}_{\rm g}\sim 10^{-4}{\hbox{$~{}h^{3}{\rm~{}Mpc}^{-3}$}}$ but the constraint improved to a few percent by increasing the number density by an order of magnitude. On the other hand, we achieved $<1\%$ constraints on $H(z)$ and $D_{\rm A}(z)$ even with the low-number density halo sample. Our full-shape analysis demonstrated that we can simultaneously investigate the Hubble tension and the flatness of the universe. Including the nonlinearity of the matter spectrum further enables us to address the $\sigma_{8}$ (or $S_{8}$) tension without combining the CMB priors.
@@ V Summary and discussion
This paper did not consider any specific galaxy sample or survey geometry. For a more realistic cosmological forecast based on simulations, we need to generate light-cone output using multiple snapshots of $N$-body simulations (see e.g., Howlett et al. (2015); Ishikawa et al. (2024)). While the halo occupation distribution (HOD) of ELGs has been studied recently (e.g., Avila et al. (2020); Okumura et al. (2021)), that of dusty galaxies still contains several uncertainties. Thus, a cosmological forecast for dusty galaxies at high redshifts requires more effort. Such detailed investigations and analyses will be performed in our future work.
@@ Acknowledgements.
We are grateful to the maintenance and administrative staff of our institutions, whose efforts in supporting our day-to-day work environment make our scientific discoveries possible. TO acknowledges support of the Taiwan National Science and Technology Council under Grants No. NSTC 112-2112-M-001-034- and NSTC 113-2112-M-001-011-. SYP and CCC acknowledge the support of the Taiwan National Science and Technology Council (111-2112M-001-045-MY3), as well as the Academia Sinica through the Career Development Award (AS-CDA-112-M02). TN acknowledges support by JSPS KAKENHI Grant Numbers JP20H05861, JP21H01081, JP22K03634, JP24H00215 and JP24H00221. KA is supported by Fostering Joint International Research (B) under Contract No. 21KK0050 and the Japan Society for the Promotion of Science (JSPS) KAKENHI Grant No. JP24K17056. We acknowledge the use of ‘the HPC facility at ASIAA’ where part of the numerical analyses were done. This work used high-performance computing facilities operated by the Center for Informatics and Computation in Astronomy (CICA) at National Tsing Hua University. This equipment was funded by the Ministry of Education of Taiwan, the Ministry of Science and Technology of Taiwan, and National Tsing Hua University.

### v2

@@ Abstract
Introduction
Cosmological models have been tested via various observations including cosmic microwave background (CMB) Hinshaw et al. (2013); Planck Collaboration et al. (2016), gravitational lensing Abbott et al. (2018); Hikage et al. (2019), galaxy redshift surveys Ivanov et al. (2020); d’Amico et al. (2020); Kobayashi et al. (2022), Type-Ia supernovae Suzuki et al. (2012); Abbott et al. (2024), etc. Cold dark matter with a cosmological constant, the $\Lambda$CDM model, has successfully explained all of these observations so far. However, there exist discrepancies between some parameters determined from the early and local universe, known as the Hubble tension (Verde et al., 2019) and the $S_{8}$ tension Abdalla et al. (2022). The Dark Energy Spectroscopic Instrument (DESI) DESI Collaboration et al. (2024), the latest galaxy survey, reported cosmological constraints that support dynamical dark energy with $w>-1$. However, combining them with other datasets still includes the cosmological constant ($w=-1$) within $2\sigma$ in a large cosmological parameter space Choudhury and Okumura (2024). Although the geometry of the universe has been found flat by the CMB, the flatness of the universe has not been constrained in late universe observations (Takada and Doré, 2015; Terasawa et al., 2022; Anselmi et al., 2023; Stevens et al., 2023; Terasawa et al., 2024). One way to tackle these issues is to extend the observations of cosmological objects to the more distant universe Okumura et al. (2016). Cosmological spectroscopic experiments at $z\gtrsim 2$ have the potential to put meaningful constraints on curvature (Takada and Doré, 2015), neutrino mass (Takada et al., 2006; Dvorkin et al., 2019), inflation (Takada et al., 2006), and dark energy (Ferraro and Wilson, 2019).
@@ Introduction
$N$-body Simulations and halo catalogs
We use $N$-body simulations run as part of an extension of the DARK QUEST project (Nishimichi et al., 2019). We employ $3000^{3}$ dark matter particles of mass $m_{p}=2.584\times 10^{10}h^{-1}M_{\odot}$ in a cubic box on the side $2h^{-1}{\rm Gpc}$. We have the data set from three independent realizations and we specifically analyze clustering of dark matter halos of the snapshots at $z=2$ and $z=3$. Halos are identified using the ROCKSTAR algorithm (Behroozi et al., 2013). Their velocities and positions are determined by the average of the member particles within the innermost 10% of the subhalo radius. The halo mass, $M_{\rm h}$, is defined by a sphere with radius within which the enclosed average density is 200 times the mean matter density.
@@
Measurements of redshift-space correlation functions
@@
×$\qquad\qquad\qquad\qquad\times\left[\xi_{\ell,k}(r_{j})-\bar{\xi}_{\ell}(r_{j})\right],$ (2)
where $\xi_{\ell,k}(r_{i})$ is the measurement from the $k$-th bootstrap realization and $\bar{\xi}_{\ell}(r_{i})=N_{\rm mock}^{-1}\sum_{k=1}^{N_{\rm mock}}\xi_{\ell,k}(r_{i})$. We create fifty realizations from each simulation. Furthermore, in measuring the redshift-space density field, we regard each direction along the three axes of simulation boxes as the line of sight. Thus, we have $N_{\rm mock}=50\times 3\times 3=450$ realizations to construct the covariance matrix.
Figure 1: Multipoles of redshift-space correlation functions of halos at $z=2$ (left) and $z=3$ (right). The upper and lower rows show the results for the low- and high-mass halos, respectively. The red and blue points are the measured monopole and quadrupole moments from $N$-body simulations. The dashed curves are the model with the input cosmological parameters. The dotted curves are the best-fitting model at $60\leq r\leq 140{\hbox{$~{}h^{-1}$}{\rm~{}Mpc}}$ (see text). Errors are the squared diagonal components of the covariance matrices.
@@
Theoretical model
Here we present a theoretical model of the correlation functions to place cosmological constraints. Since theoretical models are derived in Fourier space, we first consider a model for the redshift-space power spectrum, $P(\mathbf{k})=P(k,\mu_{\bm{k}})$, with $k=|\bm{k}|$ and $\mu_{\bm{k}}$ the directional cosine. Given the underlying matter power spectrum, cosmological information is extracted from the dynamical and geometric distortions, which are probed, respectively, via RSD and Alcock-Paczynski (AP) effects (Alcock and Paczynski, 1979; Kaiser, 1987; Seo and Eisenstein, 2003; Matsubara, 2004; Okumura et al., 2008).
We use the simplest model for RSD, the linear RSD model with a constant bias $b$ (Kaiser, 1987). It is given by
@@
These multipole power spectra can be converted into multipole correlation functions by
@@
To measure the correlation function, the halo position sampled with redshift and angular position need to be converted into the comoving positions by introducing a reference cosmology. The apparent mismatch between the reference and true cosmology causes a geometric distortion in the measured correlation function, known as the AP effect Alcock and Paczynski (1979). Including this effect, the correlation functions presented above are modulated as (Xu et al., 2013)
@@
where $H(z)$ and $D_{\rm A}(z)$ are the expansion rate and angular diameter distance, respectively, and $r_{\rm s}$ the sound horizon scale at recombination time. Quantities with the subscript $\rm{fid}$ are computed using fiducial cosmological parameters (see Introduction). Detecting anisotropies of BAO enables us to separate the AP effect from the dynamical distortion effect.
@@
Setup for cosmological analysis
In this paper, we investigate how well one can constrain cosmological models at redshifts $z>2$ in two ways: (i) constrain the growth rate and expansion rate using dynamical and geometric distortions, (ii) constrain cosmological parameters directly by utilizing the cosmological dependence of the underlying power spectrum shape.
We perform the likelihood analysis using the measured correlation functions of halos with the theoretical models described above. The $\chi^{2}$ statistic is given by
@@
Dynamical and geometric constraints
@@
Full-shape constraints
Here we utilize the cosmological information encoded in the underlying matter power spectrum and directly constrain cosmological parameters. We consider a nonflat $\Lambda$CDM model with six parameters in total, $\theta=(b,~{}\Omega_{\rm m},~{}\Omega_{\rm k},~{}H_{0},~{}A_{\rm s},~{}k_{*})$, where $A_{\rm s}$ is the amplitude of the primordial curvature perturbations. We fix the baryon density, $\omega_{\rm b}\equiv\Omega_{\rm b}h^{2}$, to the fiducial value because it is tightly constrained by CMB and big-bang nucleosynthesis Aver et al. (2013); Cooke et al. (2014); Planck Collaboration et al. (2016). We further fix the dark energy equation-of-state to a cosmological constant, $w_{0}=-1$, as the high-$z$ clustering is not sensitive to $w$. On the other hand, the high-redshift universe is a useful laboratory to test the flatness of the universe, $\Omega_{\rm k}$ (Takada and Doré, 2015).
@@
Fig. 4 shows the two-dimensional constraints on pairs of cosmological parameters. Table 3 summarizes the one-dimensional marginalized constraints. The best-fitting value of $b$ obtained here is fully consistent with that from dynamical and geometric constraints. We obtain $1.6\%$ and $2.3\%$ constraints on $\Omega_{\rm m}$ from the clustering of halos at $z=2$ and $z=3$, respectively. Similarly, we achieve constraints of $1.1\%$ and $1.3\%$ in $H_{0}$. The reported discrepancy between the values of $H_{0}$ measured using early and late universe probes is about $5-10\%$ Verde et al. (2019). Our result thus demonstrates that achieving the number density of $\bar{n}_{\rm g}\simeq 10^{-4}({\hbox{$~{}h^{3}{\rm~{}Mpc}^{-3}$}})$ over the large volume at $z\geq 2$ provides a powerful test of the Hubble tension, independent of the local universe ($z\simeq 0$) and early universe ($z\simeq 1100)$ measurements.
Note that there are certain offsets for the constraint on $\Omega_{\rm m}$ from the $z=2$ sample, though $\simeq 2\sigma$, while the results shown in the previous subsection are consistent with the input within the $1\sigma$ level for these samples. These shifts are driven by the inaccuracy of our linear theory model with the simple BAO damping on small scales, where cosmological information can still be extracted. In fact, Fig. 1 shows that the best-fit model fails to capture behavior on smaller scales than the BAO scale, especially at $z=2$, while it agrees with the measurements up to the BAO scale. This issue can be easily restored once one employs a more sophisticated model, and our main conclusion of constraining cosmology from high redshifts is unchanged.
@@
Summary and discussion
In this paper, we have studied the cosmological constraints expected from upcoming high-redshift galaxy surveys utilizing high-resolution $N$-body simulations at $z=2$ and 3. To obtain general conclusions, we did not consider specific galaxy populations or survey geometries but rather split our halo sample into low-mass and high-mass subsamples in a cubic box of side $2{\hbox{$~{}h^{-1}$}{\rm~{}Gpc}}$. We chose a conservative fitting range where a simple linear theory prediction recovered the input cosmological models. We considered two analysis methods, a template fitting for dynamical and geometric constraints and a full-shape analysis. With the former, we obtained a $\sim 10\%$ constraint on $f(z)$ with the massive halo sample with $\bar{n}_{\rm g}\sim 10^{-4}{\hbox{$~{}h^{3}{\rm~{}Mpc}^{-3}$}}$ but the constraint improved to a few percent by increasing the number density by an order of magnitude. On the other hand, we achieved $<1\%$ constraints on $H(z)$ and $D_{\rm A}(z)$ even with the low-number density halo sample. Our full-shape analysis demonstrated that we can simultaneously investigate the Hubble tension and the flatness of the universe. Including the nonlinearity of the matter spectrum further enables us to address the $\sigma_{8}$ (or $S_{8}$) tension without combining the CMB priors.
@@
This paper did not consider any specific galaxy sample or survey geometry. For a more realistic cosmological forecast based on simulations, we need to generate light-cone output using multiple snapshots of $N$-body simulations (e.g., Howlett et al. (2015); Ishikawa et al. (2024)). Although the halo occupation distribution (HOD) of ELGs has been studied recently (e.g., Avila et al. (2020); Okumura et al. (2021)), that of dusty galaxies still contains several uncertainties. Thus, a cosmological forecast for dusty galaxies at high redshifts requires more effort. Such detailed investigations and analyses will be performed in our future work.
@@ Acknowledgements.
We are grateful to the maintenance and administrative staff of our institutions, whose efforts in supporting our day-to-day work environment make our scientific discoveries possible. TO acknowledges support of the Taiwan National Science and Technology Council under Grants No. NSTC 112-2112-M-001-034- and NSTC 113-2112-M-001-011-. SYP and CCC acknowledge the support of the Taiwan National Science and Technology Council (111-2112M-001-045-MY3), as well as the Academia Sinica through the Career Development Award (AS-CDA-112-M02). TN acknowledges support by JSPS KAKENHI Grant Numbers JP20H05861, JP21H01081, JP22K03634, JP24H00215 and JP24H00221. KA is supported by Fostering Joint International Research (B) under Contract No. 21KK0050 and the Japan Society for the Promotion of Science (JSPS) KAKENHI Grant No. JP24K17056. We acknowledge ‘the HPC facility at ASIAA’ where part of the numerical analyses were done. This work used high-performance computing facilities operated by the Center for Informatics and Computation in Astronomy (CICA) at National Tsing Hua University. This equipment was funded by the Ministry of Education of Taiwan, the Ministry of Science and Technology of Taiwan, and National Tsing Hua University.